Three fair dice are tossed once. What is the probability that the faces show three consecutive integers?
Answer by stanbon(75887)

Three fair dice are tossed once. What is the probability that the faces show three consecutive integers?
----
# of ways to succeed:
There are 3 consecutive sequences possible.
There are 3*2*1 = 6 ways to achieve one of those sequences with the 3 dice.
----
# of ways to succeed = 3*6 = 18
====
# of possible outcomes when you toss 3 dice = 6^3 = 216
---
P(3 consecutive integers) = 18/216
===================
cheers,
Stan H.
===============